A thin ring of radius R and mass M rolls without s
Answer #1

here,

initial linear velocity, u = 3.50 m/s

moment of inertia of ring, I =

Part a:
linear veloicty = angular velocity * radius

angular velocity, w = u/r

option b is correct

Part b:
Total kinetic energy = 0.5*m*u^2 + 0.5*I*w^2
Total kinetic energy = 0.5*m*u^2 + 0.5*mr^2 *(u^2/r^2)
Total kinetic energy = 0.5*m(u^2 + u^2)
Total kinetic energy = mu^2

option d correct

Part c:
From conservation of energy :
Total Kinetic energy = Potential energy gained by ring
m*u^2 = m*g*h

height, h = u^2/g
height, h = 3.50^2/9.81
height, h = 1.249 m